OpenStudy (anonymous):

log3^(x-1)+log3^(x+25)=3

OpenStudy (anonymous):

is this log base 3 or log base ten?

OpenStudy (anonymous):

base is 3

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

so it is \[log_3(x-1)+log_3(x+25)=3\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry i mistyped

OpenStudy (anonymous):

if so then combine in to one logarithm using the property of logs: \[log(A)+log(B) = log(AB)\] \[log_3(x-1)(x+25)=3\]

OpenStudy (anonymous):

then write in equivalent exponential form to get \[(x-1)(x+25)=3^3=27\] \[x^2+24x-25=27\] \[x^2+24x-52=0\] \[(x-2)(x+26)=0\] \[x=2\] \[x=-26\] but -26 not a solution because you cannot take the log of a negative number, so answer should be 2

OpenStudy (anonymous):

let me know if any step is not clear.

OpenStudy (anonymous):

its clear I was just wondering why i got the -26 wrong, thanks for explaining why

OpenStudy (anonymous):

welcome.